OpenStudy (onepieceftw):

I forgot how to complete the square, how would I convert 3x^2 + 12x + 11 to vertex form?

OpenStudy (phi):

write it as \[ (3x^2+12x) + 11 \\ 3(x^2 + 4x) +11\] now inside the parens add and subtract (4/2)^2

OpenStudy (xmissalycatx):

Hey there! :) Here is an image to help you better understand of to CTS.

OpenStudy (onepieceftw):

Okay @phi , then what? This is when I factor out the trinomial x^2+4x+4? then distribute the 3

OpenStudy (phi):

you should have this: \[ y = 3x^2 + 12x + 11 \\ y= 3(x^2 +4x) + 11\] then we find (4/2)^2 = 2^2 =4 we add and subtract 4: +4-4 (which is zero, so the value is not changed) \[ y= 3(x^2 +4x+4 -4) + 11\] now group the first three terms which form a perfect square \[ x^2 +4x+4 = (x+2)^2 \] so you have \[ y= 3( (x+2)^2 -4) + 11\]

OpenStudy (phi):

now distribute the 3 and combine the numbers

OpenStudy (onepieceftw):

So I have \[y = 3(x+2)^2-1\]

OpenStudy (phi):

yes, and if we write it to match y = a(x-h)^2 + k: \[ y= 3(x- -2)^2 -1\] the vertex is -2, -1

OpenStudy (onepieceftw):

Thanks, really appreciate your help!